### v1

@@ 3 Data and Methodology
In this section, we introduce the data and explain the methodology employed in our analysis. The next subsection outlines the construction of our data set and details how we gauge players’ revealed preferences for characters. We then provide a formal review of the synthetic control estimator employed to isolate the effects of the coming-out event on players’ preferences for Graves.
@@ 3.1 Data
We obtain our data by accessing the Riot Games API, which provides us with valuable information about League of Legends matches. The game operates on multiple servers located worldwide, and we focus on specific servers for our analysis. These servers include Brazil, North and East Europe, West Europe, Korea, North Latin America, South Latin America, and North America.
@@
Within these servers, we target the top tier of the League of Legends ranked system, which comprises the top 200 or 300 players (approximately the top 0.01% of players) on each server. By targeting this specific group of players, we aim to minimize the noise that may arise from players who are not fully engaged in the game, thus reducing the risk of attenuation bias. Furthermore, this focus increases the chances that players in our data set are aware of the coming-out event, thereby strengthening the credibility of our identification strategy.151515 We acknowledge that some of these players may have followers or engage in streaming activities, which could reduce the level of anonymity within our sample. Essentially, our focus trades off a degree of anonymity for reduced attenuation bias and strengthened credibility of identification.
Specifically, we identified the players within the top tier of each server as of July 2022. For each of these players, we gathered data on all the matches they engaged in the ranked mode during the period January-July $2022$.161616 Players can choose between draft or ranked matches. Both game modes share the same mechanics and objectives. However, while draft matches are more casual and do not have consequences for players’ rankings or ratings, in ranked matches players earn or lose points based on the outcome of the match to determine their position within the ranked system. By focusing on ranked matches we aim to further reduce the risk of attenuation bias. This approach inherently includes players who were not in the top tier as of July $2022$ but have been matched with our focal players during the time span we consider.
From these data, we construct a balanced daily panel data set that tracks the usage of each character over time. We then exclude characters that have been released after the coming-out event. This results in a final data set composed of $161,756$ players, engaging in a total of $142,856$ matches played over $193$ days, encompassing a total of $159$ characters.
To gauge players’ revealed preferences for characters, we construct a metric called pick rate, which measures the frequency with which players choose a specific character in their games each day. Our primary objective is to investigate whether the disclosure of Graves’ sexual orientation influences the pick rate of this character.
@@ 3.2 Methodology
Formally, our data set comprises $n=159$ characters ($i=1,\dots,n$) observed over $T=191$ days ($t=1,\dots,T$). $T^{pre}$ is the length of the period (150 days) before the coming-out event, which occurs at time $T^{pre}+1$ (i.e., June $1^{st}$, $2022$). For each unit $i$ and time $t$, we denote the observed pick rate as $Y_{i,t}$. We represent the coming out as a binary variable $C_{i}\in\left\{0,1\right\}$ equal to one if character $i$ discloses his sexual orientation at time $T^{pre}+1$. We then posit the existence of two potential pick rates $Y_{i,t}^{c}$, where one denotes the pick rate in the absence of disclosure ($Y_{i,t}^{0}$) and the other denotes the pick rate in the presence of disclosure ($Y_{i,t}^{1}$).171717 These potential outcomes are based on Rubin’s model for causal inference [141].
@@
Since Graves’ sexual orientation has been disclosed after period $T^{pre}$, under a standard SUTVA assumption [[, see, e.g.,]]imbens2015causal we observe $Y_{1,t}=Y_{1,t}^{1}$ for all $t>T^{pre}$. Thus, as shown in equation (3.1), the challenge in estimating our causal effects of interest is to estimate $Y_{1,t}^{0}$ for $t>T^{pre}$, i.e., how Graves’ pick rates would have evolved in the absence of the disclosure. To this end, we can construct a synthetic control unit that approximates the pick rates of Graves before the coming out. The idea is that if the synthetic control and Graves behave similarly before the disclosure, then the synthetic control can serve as a valid counterfactual.
@@
$\begin{gathered}\hat{\omega}=\operatorname*{arg\,min}_{\omega\in\Omega}\ell\left(\omega\right)\\[4.30554pt] \ell\left(\omega\right)=\sum_{t=1}^{T^{pre}}\left(\sum_{i=2}^{n}\omega_{i}Y_{i,t}-Y_{1,t}\right)^{2}+\zeta^{2}T^{pre}\|\omega\|_{2}^{2},\quad\Omega=\left\{\omega\in\operatorname{\mathbb{R}}_{+}^{n-1}:\sum_{i=2}^{n}\omega_{i}=1\right\}\end{gathered}$ (3.2)
@@
$\hat{\tau}=\frac{1}{T-T^{pre}}\sum_{t=T^{pre}+1}^{T}\hat{\tau}_{t}$ (3.5)
We employ the “placebo approach” of [184] to estimate the variance of $\hat{\tau}$. We then use the estimated variance to construct asymptotically valid conventional confidence intervals.181818 The validity of this placebo approach hinges on a homoskedasticity assumption which requires that treated and control units have the same noise distribution. In general, with only one treated unit, nonparametric variance estimation for treatment effect estimators is typically impossible without a homoskedasticity assumption [184].
@@ 4 Results
In this section, we present our main results. First, we present our main findings and a series of robustness checks that validate the reliability of our estimates. We then explore the possibility of regional variations in attitudes toward the LGB community by replicating our analysis across different servers.
@@ 4.1 Main Results
We apply the synthetic control estimator of Section 3.2 to estimate the effects of the coming-out event on players’ revealed preferences for Graves. To mitigate the potential for spillover effects, we exclude Twisted Fate and other four characters (Diana, Leona, Nami, and Neeko) that were already members of the LGB community prior to the coming-out event from the donor pool.191919 Nevertheless, even if included in the donor pool, the estimator assigns these characters zero weight.
@@
Figure 4.1: Graves’ daily pick rates and synthetic control estimation results. The actual series is smoothed by a Nadaraya-Watson regression before applying the synthetic control estimator. The dashed vertical line denotes the day of disclosure, and the rainbow area highlights LGBT Pride Month.
@@
Figure 4.1 displays Graves’ actual and synthetic pick rate series, while the first column of Table 4.1 displays the point estimate and $95\%$ confidence interval for the average treatment effect.202020 Figure A.1 in Appendix A displays the identities and the contributions of the characters in the donor pool with non-zero estimated weights. Overall, our analysis suggests a substantial negative impact of the coming-out event on players’ preferences for Graves. Before the disclosure, the synthetic control estimator closely approximates the trajectory of Graves’ pick rates, providing support for the estimator’s ability to predict the counterfactual series. However, starting from June $1^{st}$, $2022$, the two series diverge substantially, with Graves’ pick rates consistently dropping below those of the synthetic control. This gap persists over time, extending even beyond the conclusion of LGBT Pride Month. The average effect is estimated to be around $-7$ percentage points and is statistically different from zero at the $5\%$ significance level.212121 To provide further evidence that our main estimate is not driven by chance, we reassigned the treatment to all control characters in our data set. This allowed us to generate a distribution of “placebo” effects, which serves as a basis for comparing the impact on Graves [[, see, e.g.,]]abadie2010synthetic, abadie2015comparative. We report the ratios of post- to pre-treatment root mean squared error (RMSE) in Figure B.1 in Appendix B, where Graves ranks second in terms of RMSE ratio. This implies that the impact on Graves is unusually large compared to the distribution of placebo effects, reinforcing the interpretation that our analysis provides significant evidence of a negative effect of the coming-out event. This implies a decline of $38.69\%$ of the pre-treatment average preferences for Graves.
Synthetic Controls Regularized Synthetic Controls (1) (2) (3) (4) All characters Only non-substitutes All characters Only non-substitutes Panel 1: All $\hat{\tau}$7.156 -6.954 -7.059 -6.933 $95\%$ CI [-13.826, -0.486] [-14.070, 0.162] [-11.900, -2.219] [-14.154, 0.286] N. Donors 4 4 4 5 RMSE 2.250 2.354 2.267 2.272 Panel 2: Europe $\hat{\tau}$8.961 -11.443 -8.368 -10.298 $95\%$ CI [-15.013, -2.910] [-19.246, -3.639] [-14.357, -2.380] [-17.566, -3.030] N. Donors 6 5 7 8 RMSE 2.151 2.524 2.189 2.477 Panel 3: Korea $\hat{\tau}$10.158 -8.983 -10.377 -8.798 $95\%$ CI [-18.022, -2.293] [-19.356, 1.391] [-18.973, -1.780] [-17.820, 0.225] N. Donors 3 3 3 3 RMSE 6.145 6.487 6.178 6.500 Panel 4: Latin America $\hat{\tau}$6.826 -6.803 -5.769 -5.456 $95\%$ CI [-11.595, -2.057] [-12.704, -0.902] [-10.509, -1.028] [-11.885, 0.973] N. Donors 4 4 7 7 RMSE 2.400 2.520 2.401 2.480 Panel 5: North America $\hat{\tau}$ 0.945 0.945 0.984 0.983 $95\%$ CI [ -4.855, 6.746] [ -6.215, 8.106] [ -4.207, 6.175] [ -6.470, 8.437] N. Donors 3 3 4 4 RMSE 4.521 4.521 4.523 4.526
Table 4.1: Point estimates and $95\%$ confidence intervals for $\hat{\tau}$. Additionally, the number of donors receiving a non-zero weight and the pre-treatment root mean squared error are displayed. The first panel reports the results obtained using all the observed matches. The remaining four panels report the results obtained using only matches from a particular region. Each column corresponds to a different specification, with the specifications differing solely in the employed estimator and donor pool composition.
We examine the robustness of our results to the choice of the estimator and the composition of units in the donor pool. In particular, we repeat our analysis employing a regularized synthetic control estimator (see Section 3.2) and we explore different donor pool configurations focusing on characters from distinct roles. Notably, Graves is predominantly designed for and played in three of the possible roles within a team. Consequently, there is a possibility of spillover effects on other characters mainly played in these positions, as players transitioning away from Graves are likely to switch to these alternatives.222222 The findings of Section 5.2 support this intuition. To mitigate this potential for spillover effects, we restrict our donor pool to characters that are “non-substitutes” of Graves, that is, those primarily designed for the remaining two roles.232323 Graves is predominantly designed for and played in the top lane, jungle, and mid lane positions. Therefore, we consider characters designed for and played in the bottom lane and support positions as “non-substitutes.”
The first panel of Table 4.1 displays the results. For any donor pool composition, the results are not sensitive to the choice of the regularization parameter. Point estimates are consistently negative across the considered specifications and are statistically different from zero at the $5\%$ significance level. An exception arises in the second column, where employing a standard synthetic control estimator while restricting the donor pool to bottom lane and support roles leads to a confidence interval that encompasses zero. The results are consistent also quantitatively across all specifications, with a decline ranging between $38.69\%$ and $37.49\%$ of the pre-treatment average preferences for Graves. Overall, these results support our main finding of a substantial negative impact of the coming-out event on players’ preferences for Graves.
We also assess the credibility of the synthetic control estimator by conducting a robustness check that artificially shifts the coming-out event ten days earlier. This backdating exercise allows us to evaluate the estimator’s predictive accuracy during a ten-day hold-out period [[, see e.g.,]]abadie2022synthetic. The upper panel of Figure B.2 in Appendix B presents the results of this analysis. We observe three key findings. First, the estimated effects remain qualitatively and quantitatively consistent, confirming a negative and persistent impact of the coming-out event on players’ revealed preferences for Graves. Second, the synthetic control estimator demonstrates a good fit during the hold-out period, indicating its ability to accurately capture Graves’ behavior prior to the disclosure. Third, the actual and the synthetic series begin to diverge on the true day of disclosure, even when the estimator has no knowledge of the actual disclosure date. The absence of estimated effects before the coming-out event also lends support to the plausibility of a no-anticipation assumption [[, see e.g.,]]abadie2021using.
Finally, we conduct an additional robustness test by performing a leave-one-out exercise, where we repeatedly estimate the synthetic control series by excluding one character with non-zero estimated weights at a time from the donor pool [[, see e.g.,]]abadie2021using. The lower panel of Figure B.2 in Appendix B presents the results of this analysis. Overall, our finding of a negative and persistent impact of the coming-out event on players’ preferences for Graves is robust to the exclusion of any particular character. Most of the leave-one-out synthetic series closely align with the main estimate, thus reinforcing the robustness of the main conclusion of our study. One leave-one-out series falls beneath the other synthetic series, suggesting a somewhat reduced, although still negative, impact. However, this series diverges from the actual series in the weeks prior to the treatment, which undermines the reliability of its results.242424 This series is obtained by excluding the character Ezreal from the donor pool. Figure A.1 shows that this character receives the largest weight in our main specification. Therefore, the divergence of this leave-one-out series from the actual series is unsurprising.

### v3

@@ 3 Data and Methodology
In this section, we introduce the data and explain the methodology employed in our analysis. The next subsection outlines the construction of our data sets and details how we gauge players’ revealed preferences for characters. We then provide a formal review of the synthetic control estimators employed to isolate the effects of the coming-out event on players’ preferences for Graves.
@@ 3.1 Data
We obtain our data by accessing the Riot Games API, which provides us with valuable information about League of Legends matches.
@@
The game operates on multiple servers located worldwide. Within these servers, we target the top tier of the League of Legends ranked system. By targeting this specific group of players, we aim to minimize the noise that may arise from players who are not fully engaged in the game, thus reducing the risk of attenuation bias.171717 Players can choose between draft or ranked matches. Both game modes share the same mechanics and objectives. However, while draft matches are more casual and do not have consequences for players’ rankings or ratings, in ranked matches players earn or lose points based on the outcome of the match to determine their position within the ranked system. By focusing on ranked matches we aim to further reduce the risk of attenuation bias. Furthermore, this focus increases the chances that players in our data set are aware of the coming-out event, thereby strengthening the credibility of our identification strategy.181818 We acknowledge that some of these players may have followers or engage in streaming activities, which could reduce the level of anonymity within our sample. Essentially, our focus trades off a degree of anonymity for reduced attenuation bias and strengthened credibility of identification.
Specifically, we identify the players within the top tier of each server as of July 2022. For each of these players, we gather raw data on the full history of matches they engaged in during the period January-July $2022$. This approach inherently includes, to a large extent, players who were not in the top tier as of July $2022$ but have been matched with our focal players during the time span we consider.
To clean the raw data, we apply a multi-step procedure. First, we aggregate server-specific regions into four primary macro-regions—Latin America, North America, Europe, and Korea—and exclude servers where match coverage is sparse, namely Japan, Oceania, Russia, and Turkey.191919 The Latin America macro-region includes servers from North Latin America, South Latin America, and Brazil. Europe combines the North/East Europe and West/South Europe servers. Korea and North America are standalone servers. Further information on server structure and geographic coverage is available at https://wiki.leagueoflegends.com/en-us/Servers. Second, we eliminate duplicate observations, that arise when the same match appears multiple times due to its inclusion in the histories of multiple players. We also drop corrupted matches, defined as those with fewer than ten unique participants, a duration exceeding two hours, or missing data. Finally, we assign each character “main” and “auxiliary” roles based on historical usage patterns and exclude matches in which characters were used in atypical roles.202020 Characters in League of Legends are explicitly designed to fulfill a primary and a secondary role. We use observed historical play patterns to empirically assign each character their main and auxiliary roles. Excluding matches where characters are used outside these roles ensures that our measure of character usage reflects meaningful player choices, rather than experimental or off-meta behavior that could introduce noise into the analysis. After these steps, our clean sample comprises 136,399136{,}399 unique players participating in 146,451146{,}451 matches.
From the clean data, we construct a character-level balanced panel that tracks the daily usage of each character across all major regions. Our final sample includes $161$ characters tracked over $194$ days. To gauge players’ revealed preferences for characters, we construct a metric called pick rate, which measures the frequency with which players choose a specific character in their games each day in each region. Our primary objective is to investigate whether the disclosure of Graves’ sexual orientation influences the pick rate of this character. We further compute daily win rates, measuring the fraction of matches in which each character is on the winning team relative to the total number of matches that character participates in on a given day. This serves as a proxy for character’s strength, allowing us to test whether shifts in character popularity are driven by changes in effectiveness rather than preferences alone.
We further construct an additional player-level daily panel that captures individual behavior over time. To ensure meaningful variation, we restrict the sample to players who participated in at least 10 matches both before and after the coming-out event, resulting in 7,4217{,}421 unique players. For each player, we compute the number of matches played and hours spent in-game each day, along with their daily win rates.212121 Hours spent in-game capture only the duration of matches. This substantially underestimates total time spent on the game, as it excludes time devoted to matchmaking, character selection, and other pre- and post-match activities. We also record the frequency with which players select specific characters and roles. This data set enables us to examine whether individual behavior adjusts in response to the coming-out event.
Table 1: Pre- and post-treatment summary statistics.
Pre-treatment Post-treatment Mean difference
Graves metrics
Graves pick rate (%) 18.56 12.32 -6.24
(3.81) (3.13)
Graves win rate (%) 50.19 49.04 -1.15
(4.82) (4.21)
Player activity
Win rate (%) 51.51 50.47 -1.04
(42.08) (40.47)
N. daily matches 2.03 2.27 +0.24
(1.62) (1.75)
Time spent in-game (hours) 0.89 1.01 +0.11
(0.72) (0.78)
Role selection patterns
Top (%) 18.51 18.86 +0.35
(36.63) (36.67)
Jungle (%) 20.93 20.83 -0.11
(38.58) (38.09)
Mid (%) 19.68 19.18 -0.50
(36.89) (36.10)
Bottom (%) 20.13 21.21 +1.09
(38.08) (38.35)
Support (%) 20.74 19.91 -0.83
(38.84) (37.77)
Notes. Sample averages are reported for each variable, followed by standard deviations in parentheses. Graves’ pick and win rates are computed from the character-level data; remaining statistics are computed from the player-level data. Time spent in-game only captures time within matches, excluding character selection and matchmaking time. The final column reports differences in means between the pre- and post-treatment periods.
Table 1 presents descriptive statistics on Graves usage and player behavior before and after the coming-out event. Graves’ pick rate declines from $18.56\%$ to $12.32\%$—a drop of more than $30\%$—suggesting a substantial shift in player preference for Graves.222222 Table A.I in Appendix A shows that, conditional on Graves’s primary role, it ranks as the second most frequently selected character before the disclosure. However, after the treatment, its ranking drops to fifth place. This decline is not accompanied by a deterioration in Graves’ win rate, which remains essentially flat. We further see no evidence that players disengaged from the game following Graves’ disclosure. On the contrary, both the average number of daily matches and hours spent in-game slightly increase post-treatment, ruling out attrition or general disengagement.232323 Figure A.I in Appendix A illustrates the distribution of players’ total and average daily activity, measured in both number of matches played and hours spent in-game. Finally, players’ role selection patterns remain largely stable across the pre- and post-treatment periods. The lack of major shifts suggests that players who stopped using Graves likely substituted toward other characters within the same role (jungle). We investigate the possibility of intra-role spillovers in greater detail in our robustness checks and find that our main results are unaffected.
@@ 3.2 Methodology
Formally, our data set comprises $n=161$ characters ($i=1,\dots,n$) observed over $T=194$ days ($t=1,\dots,T$). $T^{pre}$ is the length of the period (150 days) before the coming-out event, which occurs at time $T^{pre}+1$ (i.e., June $1^{st}$, $2022$). For each unit $i$ and time $t$, we denote the observed pick rate as $Y_{i,t}$. We represent the coming out as a binary variable $C_{i}\in\left\{0,1\right\}$ equal to one if character $i$ discloses his sexual orientation at time $T^{pre}+1$. We then posit the existence of two potential pick rates $Y_{i,t}^{c}$, where one denotes the pick rate in the absence of disclosure ($Y_{i,t}^{0}$) and the other denotes the pick rate in the presence of disclosure ($Y_{i,t}^{1}$).242424 These potential outcomes are based on Rubin’s model for causal inference [139].
@@
Since Graves’ sexual orientation has been disclosed after period $T^{pre}$, under a standard SUTVA assumption [[, see, e.g.,]]imbens2015causal we observe $Y_{1,t}=Y_{1,t}^{1}$ for all $t>T^{pre}$. Thus, as shown in equation (1), the challenge in estimating our causal effects of interest is to estimate $Y_{1,t}^{0}$ for $t>T^{pre}$, i.e., how Graves’ pick rates would have evolved in the absence of the disclosure. To this end, we can construct a synthetic control unit that approximates the pick rates of Graves before the coming out. The idea is that if the synthetic control and Graves behave similarly before the disclosure, then the synthetic control can serve as a valid counterfactual.
@@
$\begin{gathered}\hat{\omega}=\operatorname*{arg\,min}_{\omega\in\Omega}\ell\left(\omega\right),\\[4.30554pt] \ell\left(\omega\right)=\sum_{t=1}^{T^{pre}}\left(\sum_{i=2}^{n}\omega_{i}Y_{i,t}-Y_{1,t}\right)^{2}+\zeta^{2}T^{pre}\|\omega\|_{2}^{2},\quad\Omega=\left\{\omega\in\operatorname{\mathbb{R}}_{+}^{n-1}:\sum_{i=2}^{n}\omega_{i}=1\right\},\end{gathered}$ (2)
@@
$\hat{\tau}=\frac{1}{T-T^{pre}}\sum_{t=T^{pre}+1}^{T}\hat{\tau}_{t}.$ (5)
We employ the “placebo approach” of [179] to estimate the variance of $\hat{\tau}$. We then use the estimated variance to construct asymptotically valid conventional confidence intervals.252525 The validity of this placebo approach hinges on a homoskedasticity assumption which requires that treated and control units have the same noise distribution. In general, with only one treated unit, nonparametric variance estimation for treatment effect estimators is typically impossible without a homoskedasticity assumption [179].
@@ 4 Results
In this section, we present our main results. First, we present our main findings. We then investigate heterogeneity across players with different levels of in-game performance. Finally, we explore the possibility of regional variations in attitudes toward the LGB community by replicating our analysis across different macro-regions. Appendix B provides a series of robustness checks that support the reliability of our estimates.
@@ 4.1 Main Results
We apply the synthetic control estimator of Section 3.2 to estimate the effects of the coming-out event on players’ revealed preferences for Graves. To mitigate the potential for spillover effects, we exclude Twisted Fate and other four characters (Diana, Leona, Nami, and Neeko) that were already members of the LGB community prior to the coming-out event from the donor pool.262626 Nevertheless, even if included in the donor pool, the estimator assigns these characters zero weight.
@@
Figure 2: Graves’ daily pick rates and synthetic control estimation results. The actual series is smoothed by a Nadaraya-Watson regression before applying the synthetic control estimator. The dashed vertical line denotes the day of disclosure, and the rainbow area highlights LGBT Pride Month.
@@
Figure 2 displays Graves’ actual and synthetic pick rate series.272727 Figure A.II in Appendix A displays the identities and the contributions of the characters in the donor pool with non-zero estimated weights. Overall, our analysis suggests a substantial negative impact of the coming-out event on players’ preferences for Graves. Before the disclosure, the synthetic control estimator closely approximates the trajectory of Graves’ pick rates, providing support for the estimator’s ability to predict the counterfactual series. However, starting from June $1^{st}$, $2022$, the two series diverge substantially, with Graves’ pick rates consistently dropping below those of the synthetic control. This gap persists over time, extending even beyond the conclusion of LGBT Pride Month. The estimated average effect, reported in the first column of Table B.I in Appendix B, is 7.60-7.60 percentage points and is statistically different from zero at the $5\%$ significance level.282828 To provide further evidence that our main estimate is not driven by chance, we reassign the treatment to all control characters in our data set. This allows us to generate a distribution of “placebo” effects, which serves as a basis for comparing the impact on Graves [[, see, e.g.,]]abadie2010synthetic, abadie2015comparative. We report the ratios of post- to pre-treatment root mean squared error (RMSE) in Figure A.III in Appendix A, where Graves ranks third in terms of RMSE ratio. This implies that the impact on Graves is unusually large compared to the distribution of placebo effects, reinforcing the interpretation that our analysis provides significant evidence of a negative effect of the coming-out event. This implies a decline of $40.98\%$ of the pre-treatment average preferences for Graves.
We conduct a series of robustness checks, detailed in Appendix B, all of which reinforce our main finding. First, our results are robust to alternative estimation strategies and donor pool compositions: we obtain consistent estimates using a regularized synthetic control estimator and when restricting the donor pool to characters unlikely to act as substitutes for Graves. Second, we assess predictive accuracy by artificially backdating the treatment [[, see, e.g.,]]abadie2022synthetic; discrepancies between the actual and synthetic series emerge only on the true disclosure date and closely match those in the main analysis, demonstrating the estimator’s ability to forecast the counterfactual series and supporting a no-anticipation assumption. Third, a leave-one-out analysis [[, see, e.g.,]]abadie2021using confirms that no single donor character drives the results.
4.2 Player Heterogeneity
To examine whether responses to the coming-out event differ by player ability, we conduct a heterogeneity analysis based on a proxy for player skill derived from pre-treatment in-game performance. Specifically, we compute each player’s win rate prior to the treatment date, restricting the sample to the same 7,4217{,}421 users selected for our player-level data set to ensure the reliability of this metric. We then classify these players into above- and below-median skill groups based on this metric. Using this stratification, we replicate our main synthetic control analysis separately for each group. The estimation proceeds identically to our baseline specification of Section 4.1.
This stratification allows us to investigate whether the decline in Graves’ pick rate is concentrated among particular subpopulations—such as lower-skilled or more casual players—or whether it reflects a broader behavioral response. Consistent estimates across skill groups would strengthen the external validity of our findings, mitigating concerns that the observed response is driven by a narrow or non-representative segment of the player base.
Figure A.IV in Appendix A plots the distribution of overall win rates and total matches for players in each skill group. Although the win rate distributions overlap substantially, they differ in their means, with players classified as higher-skilled exhibiting slightly higher win rates on average. In contrast, the distribution of total matches is nearly identical across groups, indicating that conditioning on pre-treatment win rate does not mechanically segment players by engagement level. This is important for interpretation: it implies that our stratification holds constant (approximately) overall game activity while isolating variation in a meaningful proxy for player quality.
(a) Players below median skill
(b) Players above median skill
Figure 3: Graves’ daily pick rates and synthetic control estimation results by skill group. The actual series is smoothed by a Nadaraya-Watson regression before applying the synthetic control estimator. The dashed vertical line denotes the day of disclosure, and the rainbow area highlights LGBT Pride Month.
Figure 3 presents the results. We observe a sustained decline in Graves’ pick rate immediately following the coming-out event in both the below- and above-median skill groups. The timing and persistence of the divergence from the synthetic counterfactual are remarkably similar across the two groups and closely mirror the patterns documented in our pooled specification.292929 Appendix B shows that these findings are robust to backdating and leave-one-out exercises. These findings indicate that the estimated effect is not concentrated among players of a particular ability level but instead reflects a widespread behavioral response, thereby enhancing the external validity of our findings and mitigating concerns that the observed decline in Graves’ popularity is driven by a narrow or non-representative subset of users.303030 While qualitatively similar, the pre-treatment pick rates of Graves are lower in both skill groups relative to the main specification. This difference arises from the additional sample restriction imposed here, whereby we include only players who participated in at least 10 matches both before and after the coming-out event to ensure a meaningful proxy for skill. The main specification, by contrast, includes the full player base, potentially encompassing more casual users with higher Graves usage. As such, differences in baseline levels should be interpreted as compositional.